the supplement of an angle measures 37 degrees more than twice its complement. find the measure of an angle.
Answer by uday1435(57)

Let the angle = x
Its supplement = 180 – x
Its complement = 90 – x
Supplement measures 37 degrees more than twice its complement means
Supplement = 2 x complement + 37
(180 – x) = 2(90-x) + 37
180 – x = 180 – 2x + 37; - x + 2x = 37
Therefore x = 37.
So the answer is the angle = 37 degrees



If you still require further clarifications please write to udayakumar.t.r@gmail.com